# PillarTrack: Redesigning Pillar-based Transformer Network for Single Object Tracking on Point Clouds

### v1

@@ IV-B Modality-aware Transformer-based Backbone
Recent studies [44, 14, 45] have pointed out that the backbone designs commonly employed in image domain [46, 30, 35, 47] are sub-optimal for point cloud or depth images due to significant modality differences between RGB images and LiDAR point clouds. Particularly, backbone design in image domain tends to utilize more block layers in later stages to extract high-level semantic features for perception. For instance, ResNet-50[30] takes 6 residual blocks in the 3rd stage while the 1st stage uses 3 blocks for feature extraction. Similarly, Swin-Transformer [35] adopts 9 blocks in the 3rd stage while only 1 block is used in the 1st stage. However, we think that these design choices for distributing computations across stages were largely empirical and may not be well-suited for LiDAR point clouds. Unlike RGB images, LiDAR can capture spatial distances and shapes of objects. This means that rich and accurate geometric information of objects is already explicitly encoded in point cloud. Therefore, we assume that instead of allocating excessive computational resources to model object geometries in later stages like ResNet/Swin-Transformer in RGB images, we propose reallocating more computational resources to the early stages to effectively incorporate the geometric information carried by raw point clouds.
TABLE II: Impact of varying block quantities in the four stages of the backbone on the final results. Bold denotes the best performance.
@@
$\left[2,2,6,2\right]$ 0.29 25.7 34.9
$\left[1,1,9,1\right]$ 0.29 53.1 56.2
@@
To validate the aforementioned hypothesis for transformer-based backbone, we systematically study computation allocation by adjusting the stage compute ratio in the transformer-based backbone of SMAT [16]. It is worth noting that the setting of SMAT has the same depth number as ResNet-50, with a stage compute ratio of $[3,4,6,3]$. As shown in Table II, we used the SMAT backbone as the baseline and varied the number of depth blocks from 0 to 4 with a stride of 1 in each stage. From the results, we can find that the performance is quite sensitive to the capacity of early stage 1, while being less sensitive to the later stages (2, 3, 4). This confirmed that allocating more computational resources to the early stages is beneficial for enabling the network to better capture the geometric information carried by the raw point clouds. Based on this observation, we set the stage compute ratio of the four stages to $[3,1,1,1]$. It is worth noting that with this depth blocks setting, our modality-aware transformer backbone network achieved promising performance improvements in terms of 3.2 point in Success and 2.5 point in Precision metric, while significantly reducing the computational cost by half the GFLOPs compared to the baseline method (from 0.37 to 0.18).
Differences with Previous Study. Our results differ significantly from previous studies [44, 14], which were focused on LiDAR-based object detection tasks using CNN architecture with local receptive fields. In contrast, we focus on 3D SOT tasks within local search area and design a transformer architecture with global receptive fields. Therefore, by allocating more computational resources to the 1st stage, our network is able to capture the enough global-aware geometric information of the input point cloud. This is different with CNN-based backbone, which require more computational resources allocated to both the 1st and 2nd stages. As a result, our modality-aware transformer-based backbone design proves to be more effective.
IV-C Activation Function Selection for Point Cloud
[…]
$ReLU(x)=\max(0,x)$ (2)
$LeakyReLU(x)=\begin{cases}x&\text{if }x>0,\\ \alpha x&\text{if }x\leq 0.\end{cases}$ (3)
There is a discrepancy between the choice of activation functions natural language processing (NLP) and vision architectures. Rectified Linear Unit (ReLU)[48] is extensively used in convolutional networks (ConvNets) due to its simplicity and efficiency. With the Transformer[26] merging, Gaussian Error Linear Unit (GELU)[49] is considered as a smoother variant of ReLU and has been employed in state-of-the-art Transformer models, such as Google’s BERT [50], OpenAI’s GPT-3 [51], and Vision Transformers [35, 34], demonstrating superior performance across various language and vision tasks. However, unlike image or text data, point cloud input data contains negative values that represent corresponding geometric property. We assume that directly applying ReLU or GELU activation functions to point clouds would suppress the negative values and result in information loss. To address this issue, we explored the Leaky Rectified Linear Unit (LeakyReLU)[52], which can address the issue of neuron death and preserve more information within the range of values less than zero. As shwon in Eq. 1,2,3 The GELU and ReLU functions exhibit saturation behavior when the input values are less than zero, effectively reducing the information content. In contrast, the LeakyReLU function preserves the negative range of input values, enabling the retention of valuable information.
TABLE III: Influence of different activation functions on Car category.Bold denotes the best result
PFE Method Success Precision
GELU (SMAT[16]44footnotemark: 4) 68.9 79.6
ReLU 69.9 80.7
LeakyReLU 71.2 (+2.3) 81.8 (+2.2)
Figure 5: The impact of different activation functions on the reduction of the Epoch Loss.Inparticular, the LeakyReLU activation function has shown a more pronounced decrease in loss compared to the other two functions under consideration.
Therefore, we simply replace the GeLU activatation function in SMAT with LeakyReLU function. As shown in Tab III and Figure 5, using LeakyReLU function has better performance (+2.3 in Success and 2.2 in Precision) and lower loss value than other functions, highlighting the significant impact of activation function choice on 3D single object tracking tasks based on point cloud modality.
TABLE IV: Overall performance comparison on the KITTI dataset. The best two results are highlighted in red, blue.
Method Publication Modality Paradigm View Mean-by class Mean-by frame
Success Precision Success Precision
$M^{2}$-Track  [21] CVPR 2022 LIDAR Motion Point 61.9 80.2 60.6 80.1
SC3D[10] CVPR 2019 LIDAR Similarity Point 35.4 53.3 31.2 48.5
P2B[2] CVPR 2020 Point 39.5 53.9 42.4 60.0
3D-SiamRPN[1] IEEE Sensors J Point 43.8 58.6 46.6 64.9
MLVSNet[53] ICCV2021 Point 44.1 60.3 45.7 66.7
BAT[4] ICCV 2021 Point 47.2 65.1 51.2 72.8
PTT[3] IROS 2021 Point 48.4 63.4 55.1 74.2
LTTR[5] BMVC 2021 Voxel 50.0 67.4 48.7 65.8
V2B[22] NeurIPS 2021 Point+Voxel 52.4 65.6 58.4 75.2
PTTR[6] CVPR 2022 Point 58.4 77.8 57.9 78.1
CMT[27] ECCV 2022 Point 57.2 76.0 59.4 77.6
STNet[7] ECCV 2022 Point 63.4 81.4 61.3 80.1
C2FT[54] RAL 2022 Point 51.6 67.7 57.2 76.4
MLSET[55] RAL 2023 Point 54.2 68.9 59.6 78.4
SMAT[16] RAL 2022 Pillar 56.7 76.1 60.4 79.5
PillarTrack Ours Pillar 63.7 +7.0 78.9 +2.8 65.3 +4.9 82.2 +2.7
TABLE V: Detailed per-class performance on the KITTI dataset. The best two results are highlighted in red, blue.
Method Car-6424 Pedestrian-6088 Van-1248 Cyclist-308
Success Precision Success Precision Success Precision Success Precision
$M^{2}$-Track [21] 64.8 79.2 57.6 83.6 50.4 64.8 74.9 93.3
SC3D [10] 41.3 57.9 18.2 37.8 40.4 47.0 41.5 70.4
P2B [2] 56.2 72.8 28.7 49.6 40.8 48.4 32.1 44.7
3D-SiamRPN[1] 58.2 76.2 35.2 56.2 45.7 52.9 36.2 49.0
MLVSNet[53] 56.0 74.0 34.1 61.1 52.0 61.4 34.3 44.5
BAT [4] 60.5 77.7 42.1 70.1 52.4 67.0 33.7 45.4
PTT [3] 67.8 81.8 44.9 72.0 43.6 52.5 37.2 47.3
LTTR [5] 65.0 77.1 33.2 56.8 35.8 45.6 66.2 89.9
V2B [22] 70.5 81.3 48.3 73.5 50.1 58.0 40.8 49.7
PTTR [6] 65.2 77.4 50.9 81.6 52.5 61.8 65.1 90.5
CMT[27] 70.5 81.9 49.1 75.5 54.1 64.1 55.1 82.4
STNet [7] 72.1 84.0 49.9 77.2 58.0 70.6 73.5 93.7
C2FT [54] 67.0 80.4 48.6 75.6 53.4 66.1 38.0 48.7
MLSET [55] 69.7 81.0 50.7 80.0 55.2 64.8 41.0 49.7
SMAT [16] 71.9 82.4 52.1 81.5 41.4 53.2 61.2 87.3
PillarTrack (Ours) 74.2 +2.3 85.1 +2.7 59.7+7.6 84.7 +3.2 43.4 +2.0 51.7 77.4 +16.2 94.2 +6.9
V Experiments
KITTI Dataset. We conduct training and evaluating for our task using the KITTI [56] datasets. We follow the same setting in baseline model[16] to divide the sequences into training, validation, and testing splits, where 0-16 for training, 17-18 for validation and 19-20 for testing.
NuScenes Dataset. The NuScenes[57] dataset comprises 1000 scenes, which are allocated into 700/150/150 scenes for training, validation, and testing, respectively. Each scene is populated with approximately 300,000 points and is accompanied by comprehensive 360-degree viewpoint annotations. In our research, we adhere to the dataset configuration established by Open3DSOT[21] for the training and evaluation of our methodology. We also conduct a comparative analysis with the baseline model[16]to evaluate the efficacy and superiority of our designs.
Evaluation metrics. We use One-Pass Evaluation (OPE) [58] to measure Success and Precision. The Success measure the 3D IoU between the predicted box and the ground truth box, the Precision measures the AUC of the distance error between the center of the predicted and ground truth box.
Implementation details. We follow our baseline model[16] and employ PVTv2-b2 [37] as backbone, adhering to its original configuration. The specific settings include a head number of [1, 2, 5, 8], a depth configuration of [3, 1, 1, 1], the expansion ratio for the feed-forward layer set to [8, 8, 4, 8] and a channel number configuration of [64, 128, 320, 512]. In KITTI dataset, we set [0.1, 0.1, 4] meters as pillar size and a search area of [-3.2, -3.2, -3, 3.2, 3.2, 1] meters for the car category. In nuScenes dataset, we set [0.075, 0.075, 8] meters as pillar size and a search area of [-3.2, -3.2, -5, 3.2, 3.2, 3] meters for the car category. Additionally, the LeakyReLU activation is employed with a negative slope value of 0.01.
Training details. On the KITTI dataset, we train PillarTrack with AdamW[59] optimizer with an initial learning rate of 0.0001, a weight decay of 0.05, and a batch size of 16 on a NVIDIA 3090 GPU. The total number of training epochs was set to 72. For the NuScenes dataset, the training was conducted over 18 epochs, with a learning rate decay by a factor of 10 at epoch 14.
SMAT [16] 43.51 49.04 32.27 60.28 44.78 44.69 37.45 34.10 39.42 34.32 39.49 44.49 40.20 50.92
PillarTrack 47.12 57.72 34.18 64.93 54.82 54.41 57.70 54.63 44.68 40.73 47.70 54.48 44.59 58.86
TABLE VI: Performance comparison on the NuScenes dataset. The best two results are highlighted in red, blue.
V-A Quantitative results on KITTI.
As shown in Table IV and V, the proposed PillarTrack demonstrates significant performance improvements compared to the baseline method (SMAT [16]). Specifically, PillarTrack achieved a performance improvement of 7.0% 2.8% and 4.9% 2.7% in Success and Precision metrics in terms of mean by-class and mean-by frame setting, respectively.
In the categories of the Car, Pedestrian, Van, and Cyclist, PillarTrack outperformed SMAT by 2.3%, 7.6%, 2.0%, and 16.2% in Success, respectively. These improvements can be attributed to the effective designs of PillarTrack. Firstly, PE-PFE module could stabilize network training without information loss. Secondly, the modality-aware transformer-basde backbone focuses on capturing more geometric information from point clouds and enhance the feature representation capability of pillars. Lastly, LeakyReLU activation retains more information and help network optimization. Besides, PillarTrack also surpassed the motion-based method MMTrack [21] in the categories of Car, Pedestrian, and Cyclist, achieving state-of-the-art performance on the KITTI dataset. However, for Van, our method only has 43.4% in Success. We believe there are two main reasons for this. Firstly, the point clouds of vans are concentrated on the object surfaces, and after pillar encoding, they lack distinctive geometric features, making them more challenging to track accurately. Secondly, the limited number of training samples for the Van category may have constrained the model’s performance, as it might not have captured the full range of variations and characteristics exhibited by vans in the real world.
V-B Quantitative results on NuScenes.
As illustrated in Table VI, our PillarTrack achieves the second-best performance across the five categories on the nuScenes dataset. Compared to previous methods, M2-Tracker[21] adopts a different motion-based paradigm for object tracking, demonstrating superior results compared to the similarity-based paradigm. Meanwhile, our PillarTrack attains the best results among the similarity-based methods. Furthermore, compared to the baseline method[16], PillarTrack has demonstrated significant improvements in each category. Specifically, PillarTrack has enhanced the mean class-wise and frame-wise Success and Precision metrics by 8.21% / 9.99% and 4.39% / 7.94%, respectively. In terms of Success for each category, PillarTrack achieves notable improvements of 3.61% for Car, 1.91% for Pedestrian, 10.04% for Truck, 17.25% for Trailer, and 5.26% for Bus. Importantly, the Truck and Trailer categories demonstrate significantly enhanced performance compared to baseline. However, such improvements were not observed in the Van category in the KITTI dataset. We attribute this to two reasons: (1) The KITTI dataset only provides front-view annotations. Under the front-view perspective, the model is constrained to learn feature representations specific to the truck from a single viewpoint, which limits the acquisition of comprehensive geometric information about the Van. As a result, the Siamese network based on appearance matching struggles to capture the key geometric information of the Van. Conversely, the NuScenes dataset provides 360-degree surround view point cloud data with corresponding annotations. This enables superior learning of both appearance and geometric representations, leading to significant performance enhancement by effectively utilizing multi-view point cloud data. (2) The KITTI dataset contains a limited number of annotated van samples (1248), whereas the NuScenes dataset offers a larger sample size (13587/3352). Within our effective framework design, this facilitates richer feature representation learning, leading to performance improvements. Besides, our performance is slightly lower than that of PTTR on Car category. We think this is because PTTR utilizes both point-based and pillar-based representations to encode the input point cloud, resulting in superior performance, particularly in terms of the Success metric.

### v2

@@ IV-B Modality-aware Transformer-based Backbone
Recent studies [35, 15, 36] have pointed out that the backbone designs commonly employed in image domain [30, 32] are sub-optimal for point cloud or depth images. This is due to the significant modality differences between RGB images and LiDAR point clouds, where LiDAR inherently capture geometric semantics. Particularly, backbone design in image domain tends to utilize more computation burden in later stages to extract high-level semantic features for perception tasks. For instance, ResNet-50[30] takes 6 residual blocks in the 3rd stage while using 3 blocks in 1st stage for feature extraction. Similarly, Swin-T [32] adopts 9 blocks in the 3rd stage while only 1 block is used in the 1st stage. However, we argue that these design choices for distributing computations across stages were largely empirical and not be well-suited for LiDAR point clouds. Unlike RGB images, LiDAR can capture spatial distances and shapes of objects. This means that rich and accurate geometric information of objects is already explicitly captured in raw point cloud coordinates. Therefore, we assume that instead of allocating excessive computational resources to model object geometries in later stages like ResNet/Swin-T in RGB images, we propose reallocating more computational resources to the early stages to effectively incorporate the geometric information carried inherenty by point clouds.
To validate the aforementioned hypothesis for a transformer-based backbone, we performed t-SNE111t-SNE (t-distributed Stochastic Neighbor Embedding) is a machine learning algorithm designed for the visualization of high-dimensional data. It maps high-dimensional data to a lower-dimensional space while preserving relative distances and distribution structures, thereby facilitating exploratory data analysis and providing insights into the internal structure of the data. visualization [37] on the features processed at different stages. The effectiveness of feature extraction can be evaluated by observing cluster separation, shape, and compactness in the t-SNE visualization. Clearly separated, regular, and compact clusters suggest successful feature extraction. Initially, we visualized the features under the original blocks number setting using t-SNE, as indicated in the red box of Figure 5 (a). The visualization revealed significant overlap between clusters, irregular shapes, and lack of compactness, indicating that the features were not adequately processed under the original configuration. Therefore, keeping the total number of blocks constant, we allocated more computational resources to the earlier stages of the network. As shown in the red box of Figure 5 (b), this adjustment led to less overlap between clusters compared to the original setting, demonstrating some improvement and the learned diverse representation. However, the clusters still lacked compactness, leading us to hypothesize that there was redundant computation. To test this hypothesis, we conducted a comparative experiment with a ”3111” block configuration, as illustrated in the red box of Figure 5 (c). This setup resulted in clearly separated clusters, many of which exhibited an elliptical shape, suggesting that our hypothesis was preliminarily confirmed. In summary, the adjustments made to the allocation of computational resources across the stages of the transformer-based backbone have significantly improved the quality of feature extraction, as evidenced by the enhanced cluster separation and regularity observed in the t-SNE visualizations.
TABLE I: Performance comparison on the KITTI dataset. L means Lidar-only methods. M and S are Motion-based and Similarity-based paradigms. The best two results are highlighted in red, blue. * indicates results produced based on the official implementation.
Method Publication Paradigm View Car-6424 Pedestrian-6088 Van-1248 Cyclist-308 Mean-by class Mean-by frame
Success Precision Success Precision Success Precision Success Precision Success Precision Success Precision
$M^{2}$-Track*  [22] CVPR 2022 L+M Point 64.8 79.2 57.6 83.6 50.4 64.8 74.9 93.3 61.9 80.2 60.6 80.1
SC3D[12] CVPR 2019 L+S Point 41.3 57.9 18.2 37.8 40.4 47.0 41.5 70.4 35.4 53.3 31.2 48.5
P2B[2] CVPR 2020 Point 56.2 72.8 28.7 49.6 40.8 48.4 32.1 44.7 39.5 53.9 42.4 60.0
3D-SiamRPN[1] IEEE Sensors J Point 58.2 76.2 35.2 56.2 45.7 52.9 36.2 49.0 43.8 58.6 46.6 64.9
MLVSNet[38] ICCV2021 Point 56.0 74.0 34.1 61.1 52.0 61.4 34.3 44.5 44.1 60.3 45.7 66.7
BAT[4] ICCV 2021 Point 60.5 77.7 42.1 70.1 52.4 67.0 33.7 45.4 47.2 65.1 51.2 72.8
PTT[3] IROS 2021 Point 67.8 81.8 44.9 72.0 43.6 52.5 37.2 47.3 48.4 63.4 55.1 74.2
LTTR[5] BMVC 2021 Voxel 65.0 77.1 33.2 56.8 35.8 45.6 66.2 89.9 50.0 67.4 48.7 65.8
V2B[23] NeurIPS 2021 Point+Voxel 70.5 81.3 48.3 73.5 50.1 58.0 40.8 49.7 52.4 65.6 58.4 75.2
PTTR[6] CVPR 2022 Point 65.2 77.4 50.9 81.6 52.5 61.8 65.1 90.5 58.4 77.8 57.9 78.1
CMT[27] ECCV 2022 Point 70.5 81.9 49.1 75.5 54.1 64.1 55.1 82.4 57.2 76.0 59.4 77.6
STNet[7] ECCV 2022 Point 72.1 84.0 49.9 77.2 58.0 70.6 73.5 93.7 63.4 81.4 61.3 80.1
C2FT[39] RAL 2022 Point 67.0 80.4 48.6 75.6 53.4 66.1 38.0 48.7 51.6 67.7 57.2 76.4
MLSET[40] RAL 2023 Point 69.7 81.0 50.7 80.0 55.2 64.8 41.0 49.7 54.2 68.9 59.6 78.4
Baseline* [16] RAL 2022 Pillar 71.9 82.4 52.1 81.5 41.4 53.2 61.2 87.3 56.7 76.1 60.4 79.5
PillarTrack* Ours Pillar 74.2 85.1 59.7 84.7 61.0 69.2 78.0 95.0 68.2 83.5 66.8 83.7
Improvement +2.3 +2.7 +7.6 +3.2 +19.6 +16.0 +16.8 +7.7 +11.4 +7.2 +6.4 +4.2
To further substantiate our hypothesis from the results, we systematically study computation allocation by adjusting the stage compute ratio in the transformer-based backbone of SMAT [16]. It is worth noting that the setting of SMAT has the same depth number as ResNet-50, with a stage compute ratio of $[3,4,6,3]$. As shown in Table IV, we used the SMAT backbone as the baseline and varied the number of depth blocks from 0 to 4 with a stride of 1 in each stage. From the results, we can find that the performance is quite sensitive to the capacity of front stage, while being less sensitive to the later stages (2, 3, 4). This confirmed that allocating more computational resources to the early stages is beneficial for enabling the network to better capture the geometric information carried by the raw point clouds. Based on this observation, we set the stage compute ratio of the four stages to $[3,1,1,1]$. It is worth noting that with this depth blocks setting, our modality-aware transformer backbone network achieved promising performance improvements in terms of 3.2/in Success and 2.5/in Precision metric, while significantly reducing the computational cost by half the GFLOPs compared to the baseline method (from 0.37 to 0.18). It is noteworthy that an increase in computational load at the front end leads to a saturated performance. We surmise this is due to the saturation of the network’s parameters, where the excessive employment of parameters augments the learning burden on the network. This phenomenon naturally substantiates the existence of a performance bottleneck in the principle of computational resource allocation, illustrating that performance does not escalate linearly with additional computational resources.
TABLE II: Performance comparison on the NuScenes dataset. The best two results are highlighted in red, blue.
Baseline [16] 43.51 49.04 32.27 60.28 44.78 44.69 37.45 34.10 39.42 34.32 39.49 44.49 40.20 50.92
PillarTrack Ours 47.12 57.72 34.18 64.93 54.82 54.41 57.70 54.63 44.68 40.73 47.70 54.48 44.59 58.86
Backbone Designs Differences with Previous Study. Our results differ significantly from previous studies [35, 15], which were focused on LiDAR-based object detection tasks using CNN architecture with local receptive fields. In contrast, we focus on 3D SOT tasks within local search area and design a transformer architecture with global receptive fields. Therefore, by allocating more computational resources to the 1st stage, our modality-aware transformer-based backbone is able to capture the enough global-aware geometric semantics of the input point cloud. This is very different with CNN-based backbone, which require more computation overhead allocated to both the 1st, 2nd and 3rd stages.
V Experiments
Dataset and Evaluation metrics. Following previous works[12, 2, 4, 17, 22, 16]. we evaluate our method on KITTI [41] and nuScenes [42] datasets. For KITTI datasets, we follow the same setting in baseline model[16] to divide the sequences into training, validation, and testing splits, where 0-16 for training, 17-18 for validation and 19-20 for testing. Notably, the nuScenes[42] is a more challenging dataset due to less channel (32-line) in LiDAR and more complex scenes, containing 1000 scenes, which are allocated into 700/150/150 scenes for training, validation, and testing, respectively. For metrics, we follow the baseline and use One-Pass Evaluation (OPE) [43] to measure Success and Precision. The Success measure the 3D IoU between the predicted box and the ground truth box, the Precision measures the AUC of the distance error between the center of the predicted and ground truth box.
Implementation details. We follow our baseline model[16] and employ PVTv2-b2 [20] as backbone. The specific settings include a head number of [1, 2, 5, 8], a depth configuration of [3, 1, 1, 1], the expansion ratio for the feed-forward layer set to [8, 8, 4, 8] and a channel number configuration of [64, 128, 320, 512]. In KITTI, we set [0.05, 0.05, 4] meters as pillar size and a search area of [-3.2, -3.2, -3, 3.2, 3.2, 1] meters for the car category. In nuScenes, we set [0.075, 0.075, 8] meters as pillar size and a search area of [-3.2, -3.2, -5, 3.2, 3.2, 3] meters for the car category. To fair comparison, other training setting is the same as our baseline model, more details refer to SMAT [16].
V-A Comparison with State-of-the-art Trackers
Quantitative results on KITTI. As shown in Table I, the proposed PillarTrack demonstrates significant performance improvements compared to the baseline method (SMAT [16]). Specifically, PillarTrack achieved a performance improvement of 11.4% 7.2% and 6.4% 4.2% in Success and Precision metrics in terms of mean by-class and mean-by frame setting, respectively. In the categories of the Car, Pedestrian, Van, and Cyclist, PillarTrack outperformed SMAT by 2.3%, 7.6%, 19.6%, and 16.8% in Success, respectively. Besides, PillarTrack also surpassed the motion-based method MMTrack [22] in all classes, achieving promising results on KITTI dataset.
Quantitative results on NuScenes. The KITTI dataset has a limited number of samples, resulting in relatively modest performance gains, which may not fully reflect the capabilities of the model. Therefore, we also conduct experiments on the large-scale nuScenes [42] dataset. NuScenes contains many complex scenes with sparser point clouds, making it a more challenging benchmark. As shown in Table II, our PillarTrack achieves the second-best performance across the five categories on the nuScenes, demonstrating comparable performance among similarity-based methods. Notably, PillarTrack exhibits significant improvements over the baseline method [16] in each category, particularly with a 10.04% and 9.72% enhancement in Success and Precision metrics for the Truck category, and a 17.25% and 20.52% improvement for the Trailer category. However, our performance in the Car category is slightly lower than that of PTTR. We attribute this to PTTR’s utilization of both point-based and pillar-based representations for encoding the input point cloud, which leads to superior performance. Nevertheless, their point-based configuration may limit the potential for further quantization acceleration in onboard devices. Notably, M2-Tracker [22] adopts a motion-based paradigm for object tracking, which results in performance that surpasses similarity-based paradigms.
TABLE III: Performance comparison of PE-PFE after noise disturbance. Baseline: original setting. Rotation: Add random rotation noise within -45° to 45°. Scale: Add random scaling from 0.8 to 1.2x for the input point clouds. Translation: Add random translation from -1.2m to 1.2m for the baseline.
PFE Method Baseline Scale Translation Rotation
Suc. Prec. Suc. Prec. Suc. Prec. Suc. Prec.
PFE (SMAT[16]) 68.9 79.6 70.9 81.5 68.5 78.5 69.6 79.9
PE-PFE (Ours) 70.8 80.3 72.9 82.4 73.3 83.0 72.0 81.5
TABLE IV: Ablation of modality-aware transofrmer-based backbone. Bold denotes the best performance.
@@ V-A Comparison with State-of-the-art Trackers
$\left[6,6,2,2\right]$ 0.38 70.2 80.5
@@
V-B Ablation Study
The robustness of Pe-PFE. As shown in Tab. III, we evaluated the performance of our PE-PFE under different noise disturbance, with a particular focus on its robustness against scale, translation, and rotation variations. The results demonstrate that PE-PFE outperforms the baseline method PFE (SMAT[16]) across all conditions. These results indicate that PE-PFE effectively maintains performance stability in the presence of various noise disturbances, showcasing strong robustness.
The Design of Modality-aware Transformer-based Backbone. We conduct experiments by arranging the number of blocks per stages from 0 to 4 with a stride of 1, while setting the number of blocks in other stages to 1, as shown in Tab. IV. The results prove our previous assumption: for pillar-based LiDAR 3D SOT tasks, we should reallocate the capacity to the early stages to better integrate the geometric information carried by the raw points, instead of allocating the capacity in the later stages like image domain. Therefore, to achieve a better performance under the constraint of low computation resource, we set the stage compute ratio to (3, 1, 1, 1).
TABLE V: Ablation experiment with Blocks Number and PFE Method in the Car category in KITTI.
Method Blocks Number PFE Method Success Precision
SMAT[16] $\left[3,4,6,3\right]$ PFE 68.9 79.6
Ours $\left[3,4,6,3\right]$ PEPFE 70.8 80.3
$\left[3,1,1,1\right]$ PFE 72.1 82.1
$\left[3,1,1,1\right]$ PEPFE 74.2 85.1
Overall Ablation. Here, we conducted an ablation study to investigate the impact of the proposed PE-PFE and modality-aware backbone network design on the overall performance, as shown in Table V. Firstly, when employing the same block configuration ($\left[3,4,6,3\right]$), the PE-PFE method achieved a success rate of 70.8% and a precision of 80.3%, surpassing the baseline method SMAT. Secondly, by adjusting the block configuration to $\left[3,1,1,1\right]$, we observed that the success rate of the PE-PFE method further increased to 74.2%, with precision rising to 85.1%. This indicates that reducing the number of blocks, in conjunction with the PE-PFE strategy, effectively enhances the model’s tracking capability.
Figure 6: Visualization results comparison of our PillarTrack and baseline method (SMAT[16]) in KITTI dataset.